The sum of an angle and twice its complement is 148 degrees. find the measure of the angle and its complement.
Answer by ewatrrr(24785)

Hi,
*Note: Sum of complementary angle is 90 degrees
.
Let x represent the measure of one angle. (90-x) would then represent the other
.
question states***
x + 2*(90-x) = 148
.
simplify and solve
x + 180 - 2x = 148
.
-x = -32
x = 32
.
x = 32 degrees , its complement = 58 degrees
.
check your answer
32 + 116 = 148